Explain the significance of Noise Margin? Calculate the Noise Margin for the hardware system following static discipline with voltage levels provided below:
VIL = 1.7 V VIH = 3.2 V VOL = 0.6 V VOH = 4.4 V
1
Expert's answer
2020-12-07T03:44:30-0500

Noise Margin - this is the value of the voltage between the inverter going from a high logic level (1) to a low logic level (0) for "N_{MH}" , and Vice versa for "N_{ML}" .

"N_{MH}=V_{OH}-V_{IH}=4.4-3.2=1.2" V,

"N_{ML}=V_{IL}-V_{OL}=1.7-0.6=1.1" V.